Print

Print


Ticket Number: HEA-NOC/20080208-2      Ticket Status: CLOSE
Ticket Type:   scheduled               Resolver:      HEA-NOC
Ticket Opened: 20080208 10:08 UTC      Problem Start: 20080212 08:00 UTC
Ticket Closed: 20080213 10:15 UTC      Problem End:   20080212 09:00 UTC

Site/line:     CIT

Problem Description:
CIT is to be moved from their current connection to our Cork & DCU access
routers
to instead route to the new CRS routers via Bluenet.


Problem Effects:
While the majority of the configuration changes will be without impact, the
actual transition from the connection to the Cork & DCU PoPs to the new
topology will cause an extremely brief outage of less than one minute. 
The CIT IP connectivity service should be considered at risk during the
entire maintenance window.

Problem History:
20080208 10:01 BN Confirmed time and date with CIT.  Preparing
configurations.

20080211 16:02 BN There was a brief outage to CIT during the configuration
phase of the transition.  All connectivity has now been restored.

20080212 08:43 BN Transition work completed successfully by 08:33. 
Traffic is now flowing via Citywest as expected.  This will be monitored
for at least twenty-four hours.

20080213 10:15 BN Traffic flows have remained stable and no issues have
been reported.  Work considered successful.

Time to Fix:
1 hour

-- 
This ticket can be monitored at http://www.hea.net/tickets/20080208-2

HEAnet Limited. Registered in Ireland, No. 275301.